A CONTROL OF GEOMETRY OF ADSORBED MOLECULES IN COATINGS STUDIED BY SURFACE ENHANCED RAMAN SCATTERING

\underline{Q. XUE}, Department of Polymer Science and Engineering, College of Chemistry and Chemical Engineering, Nanjing University, Nanjing 210093, China.